Suppose you can somehow choose two people at random who took the SAT in 2014. Scores were Normally distributed with mean and stanard deviation of 1497 and 322, respectively. What is the probability that both of them scored above a 1520? Assume that the scores of the two test takers are independent.
Answer #1

For normal distribution, P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 1497

Standard deviation = 322

P(a person scored above 1520) = P(X > 1520)

= 1 - P(X < 1520)

= 1 - P(Z < (1520 - 1497)/322)

= 1 - P(Z < 0.07)

= 1 - 0.5279

= 0.4721

P(both of them scored above 1520) = 0.47212

= 0.2229